# Indecomposable set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

### v1

@@ 2. Indecomposable cycle sets of finite multipermutation level
For the reader’s convenience, we first recall a special case of the construction of indecomposable cycle sets of multipermutation level $2$ from the following result of Jedliĉka and Pilitowska [14, Prop. 5.1, Prop. 5.7].
@@
We are now able to give a more explicit version of the above result.
@@
$\psi:(\mathbb{Z}\times\bigoplus_{i\in\mathbb{Z}}\mathbb{Z})/{\sim}\to\mathbb{Z}_{m}\times(\bigoplus_{i\in\mathbb{Z}}\mathbb{Z})/H,$
@@
$(a,x)\cdot(b,y)=(b-1,y-s\chi_{0}(b)-c_{a-b}+c_{-b}).$
Now instead of starting from $H\leqslant\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}$ and $s\in\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}/H$ we can also start with an abelian group $A$, $s\in A$ and a group homomorphism $\phi:\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}\to A$, we then let $H=\ker\phi$. As the $c_{i}$, $i\neq 0$, form a basis of $\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}$, we can freely choose the images $\phi(c_{i})\in A$ as long as $\phi(c_{i})=\phi(c_{i+m})$. If we denote $\phi(c_{i})=\Phi(i)$ we see that every such $\phi$ uniquely corresponds to a map $\Phi:\mathbb{Z}_{m}\to A$ such that $\Phi(0)=0$. Using $\phi$ to identify $\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}/H$ and $A$ we find a cycle set structure on $\mathbb{Z}_{m}\times A$ given by
$(a,x)\cdot(b,y)=(b-1,y-s\chi_{0}(b)-\Phi(a-b)+\Phi(-b)).$
@@
$r\begin{pmatrix}(a,x)\\ (b,y)\end{pmatrix}=\begin{pmatrix}(b-1,y-S\chi_{0}(b-1)-\Phi(b-1-a))\\ (a+1,x+S\chi_{0}(a)+\Phi(a-b+1))\end{pmatrix}.$
@@
One easily verifies the following result:
Lemma 2.5.
Let $\psi:X\to Y$ be a surjective homomorphism of cycle sets. Then there is a unique surjective brace homomorphism $\mathcal{G}(\psi):\mathcal{G}(X)\to\mathcal{G}(Y)$ such that $\mathcal{G}(\psi)(\sigma_{x})=\sigma_{\psi(x)}$.
For any integer $n$, we denote by $\pi(n)$ the set of prime divisors of $n$.
Lemma 2.6.
[…]
  1. (1)
    $\psi=gf$,
  2. (2)
    the induced homomorphism $\mathcal{G}(g):\mathcal{G}(\tilde{X})\to\mathcal{G}(Y)$ is an isomorphism,
  3. (3)
    $\pi(|X|/|\tilde{X}|)=\pi(|K|)$.
Proof.
Recall that under the canonical embedding, $X$ embeds as a transitive cycle base into the structure brace $G(X)$. Let the ideal $K^{\prime}$ be the inverse image of $K$ under the canonical homomorphism $G(X)\to\mathcal{G}(X)$. Then it follows from [18, Theorem 1] that the orbits of $X$ under the $\lambda$-action of $K^{\prime}$ (which are the same as under the action of $K$) form a congruence relation on $X$. Let us denote the resulting cycle set $\tilde{X}$ with factor map $f:X\twoheadrightarrow\tilde{X}$. As elements of the same orbit of $K$ get identified under $\psi$, we find a well-defined surjective homomorphism $g:\tilde{X}\twoheadrightarrow Y$ such that $\psi=gf$. Now let $I$ denote the kernel of $\mathcal{G}(f):\mathcal{G}(X)\to\mathcal{G}(\tilde{X})$. As $\mathcal{G}(\psi)=\mathcal{G}(g)\mathcal{G}(f)$, we know that $I\leqslant K$. On the other hand, for any $\sigma\in K$, its permutation of $X$ induces the identity mapping on $\tilde{X}$, hence we see that $\sigma\in I$. We conclude that $I=K$ and therefore $\mathcal{G}(g)$ is an isomorphism.
As $K$ is normal in $\mathcal{G}(X)^{\circ}$, its orbits all have the same size $|X|/|\tilde{X}|$. Clearly $\pi(|X|/|\tilde{X}|)\subseteq\pi(|K|)$. On the other hand, let $p$ be a prime not contained in $\pi(|X|/|\tilde{X}|)$ and denote $K_{p}$ the unique Sylow $p$-subgroup of $K^{\circ}$. Then for every $x\in X$, the size of its orbit is coprime to $p$ hence $K_{p}$ is contained in the stabiliser of $x$ under the action of $K$. As $K^{\circ}$ acts faithfully on $X$, we conclude that $K_{p}=0$ hence $p\notin\pi(|K|)$. ∎
Theorem 2.7.
Let $X$ be a finite indecomposable cycle set of finite multipermutation level, then $\pi(|X|)=\pi(|\mathcal{G}(X)|)$.
Proof.
We prove this by induction on $|X|$. If $|X|=1$ then $\mathcal{G}(X)$ is trivial, hence the statement holds. Now assume that $|X|>1$ and consider the canonical retraction homomorphism $\psi:X\twoheadrightarrow\mathrm{Ret}(X)$. From [4, Corollary 3.9] we know that the kernel $K$ of $\mathcal{G}(X)\to\mathcal{G}(\mathrm{Ret}(X))$ is precisely $\operatorname{Soc}(\mathcal{G}(X))$. By 2.6 we know that the orbits of $K^{\circ}$ yield a cycle set $\tilde{X}$ and because $K^{\circ}$ is abelian we get that $\pi(|X|/|\tilde{X}|)=\pi(|K|)$. As $|\tilde{X}|<|X|$ we know that $\pi(|\tilde{X}|)=\pi(|\mathcal{G}(\tilde{X})|)=\pi(|\mathcal{G}(X)/K|)$, from which we conclude that
$\pi(|X|)=\pi(|X|/|\tilde{X}|)\cup\pi(|\tilde{X}|)=\pi(|K|)\cup\pi(|\mathcal{G}(X)/K|)=\pi(|\mathcal{G}(X)|).\qed$
Corollary 2.8.
Let $X$ be an indecomposable cycle set of finite multipermutation level and size $p^{n}$ for some prime $p$, then $\mathcal{G}(X)$ is a $p$-group.
Proof.
This follows directly from 2.7. ∎
@@ 3.1. Constructing the solutions
Let $X$ be an irretractable cycle set, then $\operatorname{Soc}(\mathcal{G}(X))=0$ and $\mathcal{G}(X)^{\circ}$ is not abelian.
@@
By 1.7 and 3.2 we have a unique system of imprimitivity, for $x\in X$ we denote by $\mathcal{B}_{x}$ the block containing $x$. We denote by $\mathcal{A}=\mathcal{G}\cap\mathbb{Z}_{p}^{p}$ the abelian subgroup which fixes the blocks setwise. Note that $[\mathcal{G}^{\circ}:\mathcal{A}^{\circ}]=p$.
@@
$=\lambda_{f}(f^{\circ n-1}*x)+f*x=\lambda_{f}((n-1)\gamma)+\gamma=n\gamma$
@@
Now consider $I$ be the subgroup of $\mathcal{G}_{+}$ generated by $\{a*g\mid a\in\mathcal{A},g\in\mathcal{G}\}$, which is an ideal by [18, Corollary after Proposition 6]. Because $X$ generates $\mathcal{G}^{+}$ and * is left distributive, $I$ is the subgroup of $\mathcal{G}_{+}$ generated by $\{a*x\mid a\in\mathcal{A},x\in X\}$. By the previous discussion, $I=\{0,\gamma,\ldots,(p-1)\gamma\}$. As $|I|=p$, we find that $I\subseteq\operatorname{Fix}(\mathcal{G})$. Since $I$ is a minimal normal subgroup of the nilpotent group $\mathcal{G}^{\circ}$, we also find that $I\subseteq Z(\mathcal{G}^{\circ})$, hence 1.2 yields a contradiction with the assumption that $\operatorname{Soc}(\mathcal{G})=0$. ∎
Now that we know that $\operatorname{Fix}(\mathcal{G})\cap\mathcal{A}=0$, it follows from 1.6 that $|\operatorname{Fix}(\mathcal{G})|=p$. Also, we have a semidirect product $\mathcal{G}^{\circ}=\mathcal{A}^{\circ}\rtimes\operatorname{Fix}(\mathcal{G})^{\circ}$. We use this for the following construction: for all $g\in\mathcal{G}$, we have an equality $g\circ\operatorname{Fix}(\mathcal{G})=g+\lambda_{g}(\operatorname{Fix}(\mathcal{G}))=g+\operatorname{Fix}(\mathcal{G})$. Therefore, $\mathcal{A}$ forms a system of representatives for $\mathcal{G}^{\circ}/\operatorname{Fix}(\mathcal{G})^{\circ}$ and $\mathcal{G}^{+}/\operatorname{Fix}(\mathcal{G})^{+}$. We keep the multiplication on $\mathcal{A}$ as it is, but as $\mathcal{A}$ is not necessarily closed under $+$, we define $g\oplus h$ as the unique element in $\mathcal{A}\cap(g+h+\operatorname{Fix}(\mathcal{G}))$.
@@
Suppose that $\tau:X\to\tilde{X}$ is injective, then $\mathcal{G}(\tilde{X})^{\circ}$ is isomorphic to $\mathcal{A}^{\circ}$ and each element of $\tilde{X}$ acts differently on $\tilde{X}$ by the $\tilde{\lambda}$-action, so $\tilde{X}$ is irretractable. Hence, this contradicts 3.2. Therefore, $|\tilde{X}|=p$. ∎
@@
By 3.5, for all $x\in X$, we have $x+\operatorname{Fix}(\mathcal{G})\subseteq X$. As $\mathcal{G}^{\circ}$ leaves $\operatorname{Fix}(\mathcal{G})$ invariant under the $\lambda$-action, we deduce that the cosets $x+\operatorname{Fix}(\mathcal{G})$ form a nontrivial system of imprimitivity for the $\lambda$-action of $\mathcal{G}^{\circ}$ on $X$. The uniqueness of such a system (1.7) implies that $\mathcal{B}_{x}=x+\operatorname{Fix}(\mathcal{G})$.
Moreover, it was observed earlier that $\mathcal{A}\cap(g+\operatorname{Fix}(\mathcal{G}))$ is a singleton for each $g\in\mathcal{G}$. Together with 3.6 this gives the last part of the statement. ∎
@@
Let $X$ be an indecomposable, irretractable cycle set with $\mathcal{G}(X)$ a $p$-group, then $X$ is isomorphic to a cycle set of the form $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ with the multiplication
@@
where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a nonconstant map with $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. Vice versa, this construction always results in an indecomposable, irretractable cycle set with $\mathcal{G}(X)$ a $p$-group.
@@
Finally, note that $\mathcal{G}(X)$ clearly acts transitively on the system of blocks $\{a\}\times\mathbb{Z}_{p}$. Furthermore, for a $b\in\mathbb{Z}_{p}$ with $\Phi(b)\neq 0$, we clearly have $(0,0)\cdot(b,y)=(b,y+\Phi(b))\neq(b,y)$ which shows that the $\mathcal{G}(X)$-orbit of $(b,y)$ contains at least the block $\{b\}\times\mathbb{Z}_{p}$. This implies that $\mathcal{G}(X)$ acts transitively on $X$, hence $X$ is indecomposable. Also note that $\mathcal{G}(X)$ is contained in $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}\leqslant\operatorname{Sym}_{\mathbb{Z}_{p}\times\mathbb{Z}_{p}}$, hence $\mathcal{G}(X)$ is a $p$-group. ∎
@@
Note that, using the explicit form in 3.8, we are considering the set $X\cap\mathcal{A}=\{(a,0)\mid a\in\mathbb{Z}_{p}\}$. As $\Phi$ is nonconstant, we easily see that these elements generate the whole cycle set $X$. ∎
@@ 3.2. Getting rid of redundancy and determining automorphisms
The aim of this subsection is to determine unique representatives for the irretractable cycle sets determined in Section 3 and moreover, describe their automorphism groups.
Let $\mathcal{F}_{p}$ be the set of all nonconstant functions $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ with the property that $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. $\mathcal{F}_{p}$ is acted upon by $\mathbb{Z}_{p}^{\ast}$ via $({}^{\alpha}\Phi)(A)=\alpha^{-1}\Phi(\alpha A)$. From now on, let $\mathcal{R}_{p}$ be a fixed system of representatives for this action.
@@
By 3.2, $\mathcal{G}(X)$ is nonabelian and, by 1.7, has a unique nontrivial system of imprimitivity that consists of the blocks $\{a\}\times\mathbb{Z}_{p}$. Assume that $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ comes with two cycle set operations $\cdot,\cdot^{\prime}$ that are given by the parameters $\Phi,\Phi^{\prime}\in\mathcal{F}_{p}$ and that $\phi:(X,\cdot)\to(X,\cdot^{\prime})$ is an isomorphism. Then $\phi\mathcal{G}(X,\cdot)\phi^{-1}=\mathcal{G}(X,\cdot^{\prime})$ so in particular $\phi$ must normalize the cyclic permutation action on the blocks and thus be of the form $\phi(a,x)=(\alpha a+\beta,\pi_{a}(x))$ for some $\pi_{a}\in\operatorname{Sym}_{\mathbb{Z}_{p}}$, $\alpha\in\mathbb{Z}_{p}^{*}$, $\beta\in\mathbb{Z}_{p}$. We now calculate

### v2

@@ 2. Indecomposable retractable cycle sets of size $p^{2}$
For the reader’s convenience, we first recall a special case of the construction of indecomposable cycle sets of multipermutation level $2$ by Jedlička and Pilitowska [14, Proposition 5.1, Proposition 5.7].
@@
We are now able to give a more explicit version of the above result. For an abelian group $(A,+)$ we define $\chi_{0}:A\to\mathbb{Z}$ as
$\chi_{0}(x)$ =$=\begin{cases}1&x=0\\ 0&x\neq 0\end{cases}.$
@@
$\psi:\left(\mathbb{Z}\times\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}\right)/{\sim}\to\mathbb{Z}_{m}\times\left(\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}\right)/H,$
@@
$(a,x)\cdot(b,y)=(b-1,y-\chi_{0}(b)s-c_{a-b}+c_{-b}).$
Now instead of starting from $H\leqslant\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}$ and $s\in\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}/H$ we can also start with an abelian group $A$, $s\in A$ and a group homomorphism $\phi:\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}\to A$; we then set $H=\ker\phi$. As the $c_{i}$, $i\neq 0$, form a basis of $\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}$, we can freely choose the images $\phi(c_{i})\in A$ as long as $\phi(c_{i})=\phi(c_{i+m})$. If we denote $\phi(c_{i})=\Phi(i)$ we see that every such $\phi$ uniquely corresponds to a map $\Phi:\mathbb{Z}_{m}\to A$ such that $\Phi(0)=0$. Using $\phi$ to identify $\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}/H$ and $A$ we find a cycle set structure on $\mathbb{Z}_{m}\times A$ given by
$(a,x)\cdot(b,y)=(b-1,y-\chi_{0}(b)s-\Phi(a-b)+\Phi(-b)).$
@@
$r\begin{pmatrix}(a,x)\\ (b,y)\end{pmatrix}=\begin{pmatrix}(b-1,y-\chi_{0}(b-1)S-\Phi(b-1-a))\\ (a+1,x+\chi_{0}(a)S+\Phi(a-b+1))\end{pmatrix}.$
@@ 3.1. Constructing the solutions
Let $X$ be an irretractable cycle set such that $\mathcal{G}(X)$ is a $p$-brace, then $\operatorname{Soc}(\mathcal{G}(X))=0$ and $\mathcal{G}(X)^{\circ}$ is not abelian.
@@
By 1.7 and 3.2 we have a unique system of imprimitivity, for $x\in X$ we denote by $\mathcal{B}_{x}$ the block containing $x$. We denote by $\mathcal{A}=\mathcal{G}\cap\mathbb{Z}_{p}^{p}$ the abelian subgroup of $\mathcal{G}^{\circ}$ which fixes the blocks setwise. Note that $[\mathcal{G}^{\circ}:\mathcal{A}]=p$.
@@
$=\lambda_{f}(f^{\circ n-1}*x)+f*x=\lambda_{f}((n-1)\gamma)+\gamma=n\gamma.$
@@
Now consider $I$ be the subgroup of $\mathcal{G}^{+}$ generated by $\{a*g\mid a\in\mathcal{A},g\in\mathcal{G}\}$, which is an ideal by [18, Corollary after Proposition 6]. Because $X$ generates $\mathcal{G}^{+}$ and * is left distributive, $I$ is the subgroup of $\mathcal{G}^{+}$ generated by $\{a*x\mid a\in\mathcal{A},x\in X\}$. By the previous discussion, $I=\{0,\gamma,\ldots,(p-1)\gamma\}$. As $|I|=p$, we find that $I\subseteq\operatorname{Fix}(\mathcal{G})$. Since $I$ is a minimal normal subgroup of the nilpotent group $\mathcal{G}^{\circ}$, we also find that $I\subseteq Z(\mathcal{G}^{\circ})$, hence 1.2 yields a contradiction with the assumption that $\operatorname{Soc}(\mathcal{G})=0$. ∎
Now that we know that $\operatorname{Fix}(\mathcal{G})\cap\mathcal{A}=0$, it follows from 1.6 that $|\operatorname{Fix}(\mathcal{G})|=p$. Also, we have a semidirect product $\mathcal{G}^{\circ}=\mathcal{A}\rtimes\operatorname{Fix}(\mathcal{G})^{\circ}$. We use this for the following construction: for all $g\in\mathcal{G}$, we have an equality $g\circ\operatorname{Fix}(\mathcal{G})=g+\lambda_{g}(\operatorname{Fix}(\mathcal{G}))=g+\operatorname{Fix}(\mathcal{G})$. Therefore, $\mathcal{A}$ forms a system of representatives for $\mathcal{G}^{\circ}/\operatorname{Fix}(\mathcal{G})^{\circ}$ and $\mathcal{G}^{+}/\operatorname{Fix}(\mathcal{G})^{+}$. We keep the multiplication on $\mathcal{A}$ as it is, but as $\mathcal{A}$ is not necessarily closed under $+$, we define $g\oplus h$ as the unique element in $\mathcal{A}\cap(g+h+\operatorname{Fix}(\mathcal{G}))$.
@@
Suppose that $\tau:X\to\tilde{X}$ is injective, then $\mathcal{G}(\tilde{X})^{\circ}$ is isomorphic to $\mathcal{A}$ and each element of $\tilde{X}$ acts differently on $\tilde{X}$ by the $\tilde{\lambda}$-action, so $\tilde{X}$ is irretractable. Hence, this contradicts 3.2. Therefore, $|\tilde{X}|=p$. ∎
@@
By 3.5, for all $x\in X$, we have $x+\operatorname{Fix}(\mathcal{G})\subseteq X$. As $\mathcal{G}^{\circ}$ leaves $\operatorname{Fix}(\mathcal{G})$ invariant under the $\lambda$-action, we deduce that the cosets $x+\operatorname{Fix}(\mathcal{G})$ form a non-trivial system of imprimitivity for the $\lambda$-action of $\mathcal{G}^{\circ}$ on $X$. The uniqueness of such a system (1.7) implies that $\mathcal{B}_{x}=x+\operatorname{Fix}(\mathcal{G})$.
Moreover, it was observed earlier that $\mathcal{A}\cap(g+\operatorname{Fix}(\mathcal{G}))$ is a singleton for each $g\in\mathcal{G}$. Together with the first part of this proposition, this gives the last part of the statement. ∎
@@
Let $X$ be an indecomposable, irretractable cycle set of size $p^{2}$ with $\mathcal{G}(X)$ a $p$-group, then $X$ is isomorphic to a cycle set of the form $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ with the multiplication
@@
where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. Vice versa, this construction always results in an indecomposable, irretractable cycle set with $\mathcal{G}(X)$ a $p$-group.
@@
Finally, note that $\mathcal{G}(X)$ clearly acts transitively on the system of blocks $\{a\}\times\mathbb{Z}_{p}$. For $b\in\mathbb{Z}_{p}$ with $\Phi(b)\neq 0$, we see that $(0,0)\cdot(b,y)=(b,y+\Phi(b))\neq(b,y)$ which shows that the $\mathcal{G}(X)$-orbit of $(b,y)$ contains at least the block $\{b\}\times\mathbb{Z}_{p}$. This implies that $\mathcal{G}(X)$ acts transitively on $X$, hence $X$ is indecomposable. Also note that $\mathcal{G}(X)$ is contained in $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}\leqslant\operatorname{Sym}_{\mathbb{Z}_{p}\times\mathbb{Z}_{p}}$, hence $\mathcal{G}(X)$ is a $p$-group. ∎
@@
Note that, using the explicit form in 3.8, we are considering the set $X\cap\mathcal{A}=\{(a,0)\mid a\in\mathbb{Z}_{p}\}$. As $\Phi$ is non-constant, we easily see that these elements generate the whole cycle set $X$. ∎
@@ 3.2. Getting rid of redundancy and determining automorphisms
The aim of this subsection is to determine unique representatives for the irretractable cycle sets determined in Section 3.1 and moreover, describe their automorphism groups.
Let $\mathcal{F}_{p}$ be the set of all non-constant functions $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ with the property that $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. $\mathcal{F}_{p}$ is acted upon by $\mathbb{Z}_{p}^{\ast}$ via $({}^{\alpha}\Phi)(A)=\alpha^{-1}\Phi(\alpha A)$. From now on, let $\mathcal{R}_{p}$ be a fixed system of representatives for this action.
@@
By 3.2, $\mathcal{G}(X)^{\circ}$ is non-abelian and, by 1.7, has a unique non-trivial system of imprimitivity that consists of the blocks $\{a\}\times\mathbb{Z}_{p}$. Assume that $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ comes with two cycle set operations $\cdot,\cdot^{\prime}$ that are given by the parameters $\Phi,\Phi^{\prime}\in\mathcal{F}_{p}$ and that $\phi:(X,\cdot)\to(X,\cdot^{\prime})$ is an isomorphism. Then $\phi\mathcal{G}(X,\cdot)\phi^{-1}=\mathcal{G}(X,\cdot^{\prime})$ so in particular $\phi$ must normalize the cyclic permutation action on the blocks and thus be of the form $\phi(a,x)=(\alpha a+\beta,\pi_{a}(x))$ for some $\pi_{a}\in\operatorname{Sym}_{\mathbb{Z}_{p}}$, $\alpha\in\mathbb{Z}_{p}^{*}$, $\beta\in\mathbb{Z}_{p}$. We now calculate